OpenStudy (anonymous):

Steve set up a business making friendship bracelets. He earned $70 in January. He is projected to earn 120% more in February. How much is Steve projected to make in February? $84 $98 $154 $190

jimthompson5910 (jim_thompson5910):

\[\Large {\color{red}{120}}\% \ \text{ of } \ 70 = \left( \frac{{\color{red}{120}}}{100} \right)*70 = ???\] take that result and add that to 70 to get your answer

OpenStudy (anonymous):

84?

jimthompson5910 (jim_thompson5910):

now add that to 70

jimthompson5910 (jim_thompson5910):

hmm I could be misinterpreting this. It might be simply 1.2*70 = 84 is your answer but I'm thinking that he made 84 more dollars on top of the 70

OpenStudy (anonymous):

84 + 70 is that what you mean?

jimthompson5910 (jim_thompson5910):

so I'm not sure now

OpenStudy (anonymous):

154!!

jimthompson5910 (jim_thompson5910):

yeah it might be 154

OpenStudy (anonymous):

ok is it okay if you help me with one more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

At Priscilla's Purse Warehouse, it takes 2 over 3 of a day to complete 1 over 9 of an order of purses. At this rate, how long will it take to complete the entire order of purses? 3 and 2 over 3 days 5 days 6 days 13 and 1 over 2 days

jimthompson5910 (jim_thompson5910):

we basically have this ratio 2/3 days: 1/9 jobs multiply both sides by 9 to get 9*(2/3) days: 9*(1/9) jobs 18/3 days: 9/9 jobs 6 days: 1 job so it takes 6 days to do the job

OpenStudy (anonymous):

oh i thought it would be 5 days

OpenStudy (anonymous):

you saved me there thx :)

jimthompson5910 (jim_thompson5910):

can you show me how you got 5 days? I'm curious how you got that

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

actually i did the same thing just in a different version

OpenStudy (anonymous):

i think i got 5 by mistake when multiplying

jimthompson5910 (jim_thompson5910):

ah I gotcha